A particular human hair has a Young's modulus of 3.38×109 N/m2 and a diameter of 148 μm.

If a 250 g object is suspended by the single strand of hair that is originally 20.0 cm long, by how much Δ? hair will the hair stretch?

If the same object were hung from an aluminum wire of the same dimensions as the hair, by how much Δ?(Al) would the aluminum stretch?

If the strand of hair is modeled as a spring, what is its spring constant ?hair?
Answer #1

A)

Length, dL = FL/AY

dL = (0.25 x 9.8 x 0.20)/((pi/4) x 148 x 10^-6)^2 x 3.38 x 10^9)

dL = 8.43 x 10^-3 m

B)

dL = (0.25 x 9.8 x 0.20)/((pi/4) x 148 x 10^-6)^2 x 69 x 10^9)

dL = 4.13 x 10^-4 m

C)

F/dL = (0.25 x 9.8)/0.00843 = 290.63 N/m

Comment in case any doubt please rate my answer....